INFRARED SPECTRA OF ACETYLENE-D_{2} CLUSTERS

\underline{F. MIVEHVAR}, J. NOROOZ OLIAEE, M. DEHGHANY, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada; A.R.W. MCKEL-LAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

In an effort to observe the spectra of acetylene clusters (C_{2}H_{2})_{n} with n ≥ 4 and the very weakly bound He-C_{2}H_{2} complex, spectra of acetylene-d_{2} clusters in the region of the C_{2}D_{2} ν_{3} fundamental (∼2439 cm^{−1}) were recorded using a tunable diode laser to probe a pulsed supersonic slit jet expansion. So far, several bands below 2439 cm^{−1} have been recorded. Two can be attributed to a parallel ν_{3}^{||} and perpendicular ν_{3}^{⊥} band for the T-shaped C_{2}D_{2} dimer. The interconvesion splittings are clearly seen in these bands. There are other mystery bands whose origins are being investigated both experimentally and theoretically. In this talk I will discuss possible assignments of these bands to C_{2}D_{2} clusters.